# Do We Need Subsidiarity in Software?

### v1

@@ 6.2.1. User Level:
The overarching themes for why participants selected the user level of control were functions without data collection, expectation of privacy in design, privacy, and resisting corporate excess.
@@
Functions without data collection relates to the belief that software applications could still function well without without excessive user data collection that invades privacy, and/or participant desires for a system where they could opt out of features that require excessive data collection in favor of simpler features. As examples, P5 noted that he would be just as happy to use Youtube without autofill and Google Search with its original PageRank algorithm that only relied on website links, not any user data, to rank search results. P10 expressed that a simpler advertising model could still generate revenue, appealing to television: “for TV…you’re getting commercials that aren’t targeted to you.” This theme was mentioned by 1 participant for operating systems, 3 participants for browsers, 5 participants for email, 2 participants for search engines, 2 participants for social media, 1 participant for chat applications, and was mentioned by 6 unique participants.
@@
Expectation of privacy in design relates to the belief that certain platforms — in particular, email and chat applications, as well as private social media postings — create a user expectation that their data will be private based on the affordances and purpose of the platform. P9 explained, “I think you get a reasonable expectation that your data isn’t getting [collected] because your intent is to talk to a person.” This theme was mentioned by 3 participants for email, 1 participant for social media, 3 participants for chat applications, and was mentioned by 4 unique participants.
Privacy relates to any user desire to keep their data private, including a general desire for privacy, protecting sensitive data, and freedom of speech and from government surveillance. Privacy was mentioned by 11 participants for operating systems, 9 participants for browsers, 10 participants for email, 9 participants for search engines, 7 participants for social media, 12 participants for chat applications, and was mentioned by all 16 participants.
@@
Resisting corporate excess relates to participant desires to resist corporate overreach in data usage, including the behavioral modification that comes from targeted advertising and recommendation algorithms. As examples, P5 expressed concerns with algorithms “being able to detect when someone’s about to enter a manic episode based off of their behavior on their computer, and they’ll start flooding them with ads…and they’ll spend a ton of money” P14 expressed concerns with algorithms leading to extreme ideologies. P11 expressed concerns with social media’s data-powered addictive algorithms impacting real-life relationships: “say the company is changing your algorithm to get you hooked…you’re creating unhealthy habits being online for hours and…you’re not creating authentic relationships.” Resisting corporate excess was mentioned by 2 participants for browsers, 3 participants for search engines, 4 participants for to social media, and was mentioned by 7 unique participants.
@@ 6.2.2. Community Level:
The overarching themes for why participants selected the community level of control were IT administration and security, intellectual property and data protection, organizational authority and trust, using community machines differently, productivity, collective corporate resistance, and aligning behaviors with group goals.
IT administration and security pertains to, in a workplace or other community context, allowing system administrators control of machines for purposes such as account management, security, protection against phishing or other online safety, and compliance with software requirements. IT administration and security was mentioned by 7 participants for operating systems, 7 participants for browsers, 5 participants for email, 2 participants for search engines, 2 participants for social media, 2 participants for chat applications, and was mentioned by 10 unique participants.
@@
Intellectual property and data protection means that in a workplace context, the employer should have control over employee data for the purpose of protecting the company’s intellectual property and sensitive data. P11, who works in our university’s development office, mentioned the following as a reason why work emails should be controlled by the employer: “what type of information are we emailing donors? Or the sensitive information that we have about donors: who are we sending that to? It should be only internal.” Intellectual property and data protection was mentioned by 6 participants for operating systems, 1 participant for browsers, 4 participants for email, 2 participants for search engines, 2 participants for chat applications, and was mentioned by 9 unique participants.
@@
Organizational authority and trust pertains to a willingness to have data be controlled at the community level because of trust in those institutions or belief in their right to control our data as part of our membership. Reasons for trusting such institutions include belief in the mission of the organization, smaller organizations having different incentives than large corporations, and smaller organizations being easier to hold accountable than large corporations. P3 said this about why she trusts her employer, our university, with her data: “I agree with its values so strongly that it makes me actually have a strong desire to have control on that level, even though I’m sure there are not many people who work for a company that they feel so strongly connected in terms of values.” Participants also mentioned that our employers have a right to control our work data as part of our employment agreement, and participation in community organizations comes with a degree of buy-in or trust of the community. P2 described this in the professional context: “the act of entering into a working agreement with a company is very different than just the act of using a service like Google search. If you’re signing up to work at a company or to go to a school, I think there’s a much more reasonable degree to which they can have restrictions or have access to certain elements of your data.” Another element of this theme was a desire from three participants for community collectives to help protect the security of software applications or provide other data-driven features, such that the individual user neither has to do these things themselves nor surrender their data to a large tech company to provide these features. Organizational authority and trust was mentioned by 4 participants for operating systems, 5 participants for browsers, 4 participants for email, 4 participants for search engines, 3 participants for social media, 3 participants for chat applications, and was mentioned by 11 unique participants.
Using community machines differently encompasses participant comments that they do not use community software for personal usage (especially when that community is their workplace), thus they are comfortable with control at the community level because they are not sharing any personal data. Using community machines differently was mentioned by 3 participants for operating systems, 2 participants for browsers, 1 participant for email, 1 participant for search engines, 1 participant for chat applications, and was mentioned by 6 unique participants.
@@
Productivity relates to optimizing the use of different tools. In the workplace context this could mean getting “the sense of the patterns of work, tools that are being used” (P12) from operating system or browsing data. In the context of a community-run social media platform, this could mean an administrator having access to user data for the purpose of “tailoring recommendations based on….what the group would be interested in” (P6). Productivity was mentioned by 1 participant for browsers, 1 participant for search engines, 1 participant for social media, 1 participant for chat applications, and was mentioned by 4 unique participants.
@@
Collective corporate resistance refers to the sentiment that community organizations could protect users from corporate abuses better than an individual user could protect themselves. P9 said, “The responsibility shouldn’t go to the individual. From an individual level, it’s hard to hold the corporation accountable because you don’t have power as one person.” Collective corporate resistance was mentioned by 1 participant for operating systems, 1 participant for browsers, 1 participant for email, 1 participant for search engines, and was mentioned by 3 unique participants.
@@
Aligning behaviors with group goals refers to acceptable use of community-owned machines (e.g., as P5 said, “if they’re installing stuff they’re not supposed to be installing, like I gave [a computer] to someone for work and they downloaded Minecraft. Or, if they’re mining Bitcoin with their computer instead of using it for school”), respectful conduct on email, chat, and social media applications (e.g., no hate speech or harassment), and keeping forums on topic (e.g., as P7 said, “we’re a soccer group. We don’t want to be publishing about crafts or science”). Aligning behaviors with group goals was mentioned by 6 participants for operating systems, 8 participants for browsers, 5 participants for email, 5 participants for search engines, 8 participants for social media, 8 participants for chat applications, and was mentioned by 15 unique participants.
@@ 6.2.3. Platform Level:
The overarching themes for why participants selected the platform level of control were operational reliability, user experience, assuming less privacy in usage, and free to use.
Operational reliability pertains to a platform’s ability to automatically fix bugs and protect security. It was mentioned by 1 participant for operating systems, 1 participant for browsers, 1 participant for social media, and was mentioned by 3 unique participants.
User experience relates to a platform’s ability to improve user experience and build new features because of the large amounts of data collected. Our participants mentioned features like improved AI models and search results, spam filters, content recommendation algorithms, and trend forecasting. User experience was mentioned by 1 participant for operating systems, 2 participants for browsers, 3 participants for email, 5 participants for search engines, 4 participants for social media, 1 participant for chat applications, and was mentioned by 6 unique participants.
@@
Assuming less privacy in usage means that on certain software platforms or in certain contexts, users will not store sensitive data. Because of the actions they take personally to protect their data, they are comfortable with the platform level of control. P9 explained that he assumes anything he searches on Google could be used as training data for algorithms and thus doesn’t search for sensitive things: “I’m communicating with the machine on my Google searches. So there’s kind of a reasonable expectation that that they can train on my data.” P3, a university employee who works with sensitive student data, shared how she takes personal ownership to protect student data: “If a student sends me a transcript and it has a social security number on it, I’m gonna delete that file from my computer.” Assuming less privacy in usage was mentioned by 1 participant for operating systems, 2 participants for search engines, and was mentioned by 3 unique participants.
@@
Free to use relates to a software application being free to use because of user data collection for targeted advertising and other purposes. Free to use was mentioned by 2 participants for browsers, 2 participants for social media, and was mentioned by 3 unique participants.

### v2

@@ 6.2.1. User Level:
The overarching themes for why participants selected the user level of control were functions without data collection, expectation of privacy in design, privacy, and resisting corporate excess. The number of participants who mentioned each theme per software application type, as well as the total number of unique participants who mentioned each theme across any software application type, can be seen in Figure 17.
Figure 17. Heat map of themes participants identified to justify user-level control, broken down by each software platform type. We also note the number of unique participants who mentioned this theme across any software platform type. The privacy theme was mentioned by the most participants across all software platform types.
A heatmap. The table has five columns: “Functions Without Data Collection,” “Expectation of Privacy in Design,” “Privacy,” and “Resisting Corporate Excess” under the main heading “Theme Justifying User-Level Control,” and “Application” as the row header. The rows are “OS,” “Browser,” “Email,” “Search,” “Social Media,” “Chat,” and “Total Unique Participants.” A legend on the bottom indicates the color shading based on the number of participants: Gray: 0 participants, Light Blue: 1-3 participants, Medium-Light Blue: 4-7 participants, Blue: 8-11 participants, Dark Blue: 12-16 participants. Here is a breakdown of the data in the heatmap: OS: “Functions Without Data Collection” column is light blue with the number 1. “Expectation of privacy in design” is gray with 0. “Privacy” is blue with 11. “Resisting Corporate Excess” is gray with 0. Browser: “Functions Without Data Collection” column is light blue with the number 3. “Expectation of privacy in design” is gray with 0. “Privacy” is blue with 9. “Resisting Corporate Excess” is light blue with 2. Email: “Functions Without Data Collection” column is medium-light blue with the number 5. “Expectation of privacy in design” is light blue with 3. “Privacy” is dark blue with 10. “Resisting Corporate Excess” is gray with 0. Search: “Functions Without Data Collection” column is light blue with the number 2. “Expectation of privacy in design” is gray with 0. “Privacy” is blue with 9. “Resisting Corporate Excess” is light blue with 3. Social media: “Functions Without Data Collection” column is light blue with the number 2. “Expectation of privacy in design” is light blue with 1. “Privacy” is medium-light blue with 7. “Resisting Corporate Excess” is medium-light blue with 4. Chat: “Functions Without Data Collection” column is light blue with the number 1. “Expectation of privacy in design” is light blue with 3. “Privacy” is dark blue with 12. “Resisting Corporate Excess” is medium-light gray with 0. Total unique participants: “Functions Without Data Collection” column is medium-light blue with the number 6. “Expectation of privacy in design” is medium-light blue with 4. “Privacy” is dark blue with 16. “Resisting Corporate Excess” is gray medium-light blue with 7.
@@
Functions without data collection relates to the belief that software applications could still function well without without excessive user data collection that invades privacy, and/or participant desires for a system where they could opt out of features that require excessive data collection in favor of simpler features. As examples, P5 noted that he would be just as happy to use Youtube without autofill and Google Search with its original PageRank algorithm that only relied on website links, not any user data, to rank search results. Similarly, P10 expressed a desire for the ability to opt out of such features that improve user experience but require excessive data collection. P10 also expressed that a simpler advertising model could still generate revenue, appealing to television: “for TV…you’re getting commercials that aren’t targeted to you.” Importantly, some of our participants held that higher-level control was necessary for certain software types but user-level control was possible for others. As an example, while P2 believed that platform-level control was necessary for social media and search to keep them free, he believed that this higher-level control was not necessary for all chat platforms due to different business models employed: “iMessage is notably, in theory, more secure than a lot of others, because Apple sees it as core to their model of selling iPhones. So the value proposition is coming through something different.”
@@
Expectation of privacy in design relates to the belief that certain platforms — in particular, email and chat applications, as well as private social media postings — create an expectation that data will be private based on the affordances and purpose of the platform. P9 explained, “I think you get a reasonable expectation that your data isn’t getting [collected] because your intent is to talk to a person” and compared platforms like email and chat to platforms like search engines: “in one case, you’re communicating to…a search service….you’re communicating with the machine versus you’re communicating between two people when you’re communicating via email…If I’m communicating to the machine, whatever I put into the machine that’s on me…But in [the case of email or chat], I’m communicating to another person.” P7 expressed that in the case of social media, when a user makes a post that’s private, it creates an expectation that the post will not only be viewable only to their approved connections but also that the data from that post will not be collected by the social media platform. Finally, some participants simply expressed feelings of privacy as why their data should be kept private. P9 said about email, “I feel like communication is more private” and P10 said about chat platforms, “I just feel like it’s private when you’re messaging the other person.”
Privacy relates to any user desire to keep their data private, including a general desire for privacy, protecting sensitive data, and freedom of speech and from government surveillance. In relation to operating systems, P6 said, “Ideally, most of it should be controlled by the user…You could have so much information, like your social security number is probably on there somewhere.” P7 similarly said, “I probably have…personal pictures, information, medical records, textual notes…you just want to keep private your own personal information to avoid fraud.” In the context of email and chat platforms, participants didn’t want to “share any of my information of what I’m talking about with my friends or personal lives” (P15) and didn’t want to have “sensitive information…leaked” (P11). Other participants expressed a desire to not have their data be “sold to somewhere I don’t know…or be used for destructive things” (P11) and also did not want to be tracked. Fear of government surveillance and a desire for the “freedom to associate” (P14) was also a concern: “maybe if someday…the government had ties to certain browsing things…I might feel like I’m not allowed to look certain things up or go to certain news websites…or I might be put on some watch list for going to a certain website” (P5). Both P13 and P16 mentioned a “right to privacy,” indicating a belief that privacy is more than a mere personal preference.
@@
Resisting corporate excess relates to participant desires to resist corporate overreach in data usage, including the behavioral modification that comes from targeted advertising and recommendation algorithms. As examples, P5 expressed concerns with algorithms “being able to detect when someone’s about to enter a manic episode based off of their behavior on their computer, and they’ll start flooding them with ads…and they’ll spend a ton of money.” P14 expressed concerns with algorithms leading to extreme ideologies. P11 expressed concerns with social media’s data-powered addictive algorithms impacting real-life relationships: “say the company is changing your algorithm to get you hooked…you’re creating unhealthy habits being online for hours and…you’re not creating authentic relationships.”
@@ 6.2.2. Community Level:
The overarching themes for why participants selected the community level of control were IT administration and security, intellectual property and data protection, organizational authority and trust, using community machines differently, productivity, collective corporate resistance, and aligning behaviors with group goals. We note that while our goal is to assess why a particular level of control is necessary, the themes of organizational authority and trust and using community machines differently do not necessarily reflect the necessity of community-level control. Rather, these themes help further explain why our participants were willing to surrender control of their data to the community level in exchange for other benefits they found necessary. We report these themes although they are not explanations of necessity because they further illuminate our participant attitudes towards levels of control. The number of participants who mentioned each theme per software application type, as well as the total number of unique participants who mentioned each theme across any software application type, can be seen in figure 18.
Figure 18. Heat map of themes participants identified to justify community-level control, broken down by each software platform type. We also note the number of unique participants who mentioned this theme across any software platform type. The aligning behaviors with group goals theme was mentioned by the most participants across all software platform types. IT administration and security, intellectual property and data protection, and organizational authority and trust also had high levels of responses across our participants.
A heatmap. The table has eight columns: “IT Administration and Security,” “Intellectual Property and Data Protection,” “Organizational Authority and Trust,” “Using Community Machines Differently,” “Productivity,” “Collective Corporate Resistance,” and “Aligning Behaviors With Group Goals” under the main heading “Theme Justifying Community-Level Control,” and “Application” as the row header. The rows are “OS,” “Browser,” “Email,” “Search,” “Social Media,” “Chat,” and “Total Unique Participants.” A legend on the right indicates the color shading based on the number of participants: Dark Blue: 12-16 participants, Blue: 8-11 participants, Medium-Light Blue: 4-7 participants, Light Blue: 1-3 participants, Gray: 0 participants. Here is a breakdown of the data in the heatmap: OS: “IT Administration and Security” is medium-light blue with the number 7. “Intellectual Property and Data Protection” is medium blue with the number 6. “Organizational Authority and Trust” is medium blue with the number 4. “Using Community Machines Differently” is light blue with the number 3. “Productivity” is gray with the number 0. “Collective Corporate Resistance” is light blue with the number 1. “Aligning Behaviors With Group Goals” is medium-light blue with the number 6. Browser: ‘IT Administration and Security” is medium-light blue with the number 7. “Intellectual Property and Data Protection” is light blue with the number 1. “Organizational Authority and Trust” is medium blue with the number 5. “Using Community Machines Differently” is light blue with the number 2. “Productivity” is light blue with the number 1. “Collective Corporate Resistance” is light blue with the number 1. “Aligning Behaviors With Group Goals” is blue with the number 8. Email: ‘IT Administration and Security” is medium-light blue with the number 5. “Intellectual Property and Data Protection” is medium-light blue with the number 4. “Organizational Authority and Trust” is medium-light blue with the number 4. “Using Community Machines Differently” is light blue with the number 1. “Productivity” is gray with the number 0. “Collective Corporate Resistance” is light blue with the number 1. “Aligning Behaviors With Group Goals” is medium-light blue with the number 5. Search: ‘IT Administration and Security” is light blue with the number 2. “Intellectual Property and Data Protection” is light blue with the number 2. “Organizational Authority and Trust” is medium-light blue with the number 4. “Using Community Machines Differently” is gray with the number 0. “Productivity” is light blue with the number 1. “Collective Corporate Resistance” is gray with the number 0. “Aligning Behaviors With Group Goals” is blue with the number 8. Social media: ‘IT Administration and Security” is light blue with the number 2. “Intellectual Property and Data Protection” is gray with the number 0. “Organizational Authority and Trust” is light blue with the number 3. “Using Community Machines Differently” is gray with the number 0. “Productivity” is light blue with the number 1. “Collective Corporate Resistance” is gray with the number 0. “Aligning Behaviors With Group Goals” is blue with the number 8. Chat: ‘IT Administration and Security” is light blue with the number 2. “Intellectual Property and Data Protection” is light blue with the number 2. “Organizational Authority and Trust” is light blue with the number 3. “Using Community Machines Differently” is light blue with the number 1. “Productivity” is light blue with the number 1. “Collective Corporate Resistance” is gray with the number 0. “Aligning Behaviors With Group Goals” is blue with the number 8. Total unique participants: ‘IT Administration and Security” is blue with the number 10. “Intellectual Property and Data Protection” is blue with the number 9. “Organizational Authority and Trust” blue with the number 11. “Using Community Machines Differently” is medium-light blue with the number 6. “Productivity” is medium-light blue with the number 4. “Collective Corporate Resistance” is light blue with the number 3. “Aligning Behaviors With Group Goals” is dark blue with the number 15.
IT administration and security pertains to, in a workplace or other community context, allowing system administrators control of machines for purposes such as account management, security, protection against phishing or other online safety, and compliance with software requirements. Our participants found it necessary in professional or other group contexts to have this higher level of control to ensure compliance with group rules and to protect the security of the organization and all members of the group. P6 highlighted the importance of increased security in professional settings saying, “for a university, there might be more security restrictions, so in that case, they’re gonna have a little more control.” Increased security in community contexts can help protect all members of the community: if one member of the community “download[s] some malware or virus, [they could] spread the virus across [our university]” (P7). Community-level control can also help community members receive the technical support they need to use their machines: P6 mentioned the importance of having an IT administrator set up accounts for the virtual machines used in his research lab. Finally, there is a particular importance in having community-level security protections when minors are members of the community. On this note, P12 said, “working with minors, there’s particular sensitivity that they should have protections extended in their digital space….there’s lots of things that the quick response from a community-level awareness could could help prevent.”
@@
Intellectual property and data protection means that in a workplace context, the employer should have control over employee data for the purpose of protecting the company’s intellectual property and sensitive data. P11, who works in our university’s development office, mentioned the following as a reason why work emails should be controlled by the employer: “what type of information are we emailing donors? Or the sensitive information that we have about donors: who are we sending that to? It should be only internal.” Other data to keep private includes “proprietary information” (P6) and “copyright material…[and] intellectual property” (P14), for example if a company is “developing some sort of product that’s not yet public…and [someone is] emailing it to someone out of the organization” (P10). P8 and P14 both mentioned keeping data private to protect an organization’s image or public relations.
@@
Organizational authority and trust pertains to a willingness to have data be controlled at the community level because of trust in those institutions or belief in their right to control our data as part of our membership. Reasons for trusting such institutions include belief in the mission of the organization, smaller organizations having different incentives than large corporations, and smaller organizations being easier to hold accountable than large corporations. P3 said this about why she trusts her employer, our university, with her data: “I agree with its values so strongly that it makes me actually have a strong desire to have control on that level, even though I’m sure there are not many people who work for a company that they feel so strongly connected in terms of values.” Participants also mentioned that our employers have a right to control our work data as part of our employment agreement, and participation in community organizations comes with a degree of buy-in or trust of the community. P2 described this in the professional context: “the act of entering into a working agreement with a company is very different than just the act of using a service like Google Search. If you’re signing up to work at a company or to go to a school, I think there’s a much more reasonable degree to which they can have restrictions or have access to certain elements of your data.” Another element of this theme was a desire from three participants for community collectives to help protect the security of software applications or provide other data-driven features, such that the individual user neither has to do these things themselves nor surrender their data to a large tech company to provide these features.
Using community machines differently encompasses participant comments that they do not use community software for personal usage (especially when that community is their workplace), thus they are comfortable with control at the community level because they are not sharing any personal data. This encompasses both users changing their behavior on community-owned machines (P7 said, “If it’s the work computer, then I shouldn’t be putting too much personal information there”) and users modulating their expectations of privacy on community-owned machines (P15 said, “I don’t necessarily have a problem with them monitoring it if they’re not sharing it with anything else, because I’m just using these for professional purposes”).
@@
Productivity relates to optimizing the use of different tools. In the workplace context this could mean getting “the sense of the patterns of work, tools that are being used” (P12) from operating system or browsing data. In the context of a community-run social media platform, this could mean an administrator having access to user data for the purpose of “tailoring recommendations based on….what the group would be interested in” (P6). This can help communities run more effectively: P12 noted this necessity in the context of businesses, saying, “to see what they’re paying money for, [if] a program isn’t being used….[this] would be helpful so that they can monitor their expenses and their investments and different kinds of programs, and…the general satisfaction from employees with with what they’re offering.”
@@
Collective corporate resistance refers to the sentiment that community organizations could protect users from corporate abuses better than an individual user could protect themselves, thus there is a necessity to have data be controlled at the community level for the purposes of protecting the individual users. P9 said, “The responsibility shouldn’t go to the individual. From an individual level, it’s hard to hold the corporation accountable because you don’t have power as one person.”
@@
Aligning behaviors with group goals refers to acceptable use of community-owned machines (e.g., as P5 said, “if they’re installing stuff they’re not supposed to be installing, like I gave [a computer] to someone for work and they downloaded Minecraft. Or, if they’re mining Bitcoin with their computer instead of using it for school”), respectful conduct on email, chat, and social media applications (e.g., no hate speech or harassment), and keeping forums on topic (e.g., as P7 said, “we’re a soccer group. We don’t want to be publishing about crafts or science”). Ensuring acceptable use of machines in community contexts can be necessary for the communities to run effectively and achieve their goals.
@@ 6.2.3. Platform Level:
The overarching themes for why participants selected the platform level of control were operational reliability, user experience, assuming less privacy in usage, and free to use. We again note that while our goal is to assess why the platform level of control is necessary, assuming less privacy in usage does not necessarily reflect the necessity of platform-level control. Again, this theme helps to further explain why our participants were willing to surrender control of their data to the platform level in exchange for other benefits they found necessary. The number of participants who mentioned each theme per software application type, as well as the total number of unique participants who mentioned each theme across any software application type, can be seen in figure 19.
Figure 19. Heat map of themes participants identified to justify platform-level control, broken down by each software platform type. We also note the number of unique participants who mentioned this theme across any software platform type. User experience was mentioned by the most participants across all software platform types.
A heatmap. The table has five columns: “Operational Reliability,” “User Experience,” “Assuming Less Privacy in Usage,” and “Free to Use” under the main heading “Theme Justifying Platform-Level Control,” and “Application” as the row header. The rows are “OS,” “Browser,” “Email,” “Search,” “Social Media,” “Chat,” and “Total Unique Participants.” A legend on the bottom indicates the color shading based on the number of participants: Gray: 0 participants, Light Blue: 1-3 participants, Medium-Light Blue: 4-7 participants, Blue: 8-11 participants, Dark Blue: 12-16 participants. Here is a breakdown of the data in the heatmap: OS: “Operational Reliability” is light blue with the number 1. “User Experience” is light blue with the number 1. “Assuming Less Privacy in Usage” is light blue with the number 1. “Free to Use” is gray with 0. Browser: “Operational Reliability” is light blue with the number 1. “User Experience” is light blue with the number 2. “Assuming Less Privacy in Usage” is gray with 0. “Free to Use” is light blue with 2. Email:“Operational Reliability” is gray with 0. “User Experience” is light blue with the number 3. “Assuming Less Privacy in Usage” is gray with 0. “Free to Use” is gray with 0. Search: “Operational Reliability” is gray with 0. “User Experience” is medium-light blue with the number 5. “Assuming Less Privacy in Usage” is light blue with the number 2. “Free to Use” is gray with 0. Social media: “Operational Reliability” is light blue with the number 1. “User Experience” is medium-light blue with the number 4. “Assuming Less Privacy in Usage” is gray with 0. “Free to Use” is light blue with 2. Chat: “Operational Reliability” is gray with 0. “User Experience” is light blue with the number 1. “Assuming Less Privacy in Usage” is gray with 0. “Free to Use” is gray with 0. Total unique participants: “Operational Reliability” is light blue with the number 3. “User Experience” is medium-light blue with the number 6. “Assuming Less Privacy in Usage” is light blue with the number 3. “Free to Use” is light blue with 3.
Operational reliability pertains to a platform’s ability to automatically fix bugs and protect security. P3 highlighted the importance of data control at the platform level “so that my company’s IT services are not doing all of the bug crashes and fixes.” Additionally, P1 expressed concern that without access to data software companies cannot know what needs to be improved: “if you have no control over any data, no examples…then how can you improve your product security to prevent future attacks?” (P1).
User experience relates to a platform’s ability to improve user experience and build new features because of the large amounts of data collected. Our participants mentioned features like improved AI models and search results, spam filters, content recommendation algorithms, and trend forecasting. The participants who cited user experience as a reason for selecting platform-level control believed that these platforms would not work without access to user data to improve user experience: “getting information about what that person [is] interested in, recommendations…I think that is necessary for those apps to function and that’s why people even get them in the first place” (P6). One participant even went so far as to say that “the company…should have a right to see how [the product] is being used” (P8).
@@
Assuming less privacy in usage means that on certain software platforms or in certain contexts, users will not store sensitive data. Because of the actions they take personally to protect their data, they are comfortable with the platform level of control. P9 explained that he assumes anything he searches on Google could be used as training data for algorithms and thus doesn’t search for sensitive things: “I’m communicating with the machine on my Google searches. So there’s kind of a reasonable expectation that that they can train on my data.” P3, a university employee who works with sensitive student data, shared how she takes personal ownership to protect student data: “If a student sends me a transcript and it has a social security number on it, I’m gonna delete that file from my computer.”
@@
Free to use relates to a software application being free to use because of user data collection for targeted advertising and other purposes. Participants who mentioned this theme as a reason for selecting platform-level control found it necessary to have everyday platforms be free: “I do value having something like a web browser be free” (P3).